# Smooth Tests for Normality in ANOVA

### v3

@@ Appendix A Appendix
from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\hat{\Psi}^{2}_{1}$, $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,\hat{K}=2\right)$
$\approx$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$=$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)-\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right).$
As $N(\hat{\Psi}^{2}_{2}-\hat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
@@
$\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>x\right)$
$\leq$ $\mathrm{P}\left(N\hat{\Psi}^{2}_{2}>2\log N\right)$
@@
where we use the distribution function of the limiting $\chi^{2}_{2}$ for $N\hat{\Psi}^{2}_{2}$. Hence $\mathrm{P}(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$ converges to 0 much slower than $\mathrm{P}(N\hat{\Psi}^{2}_{2}>x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N)$, and thus the latter can be neglected. Therefore, for $x\geq 2\log N$, we have the approximation
$\mathrm{P}\left(N\hat{\Psi}^{2}_{1}\leq x,N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\leq\log N\right)+\mathrm{P}\left(N\hat{\Psi}^{2}_{2}-N\hat{\Psi}^{2}_{1}\geq\log N\right)$
@@
For $\log N<x<2\log N$, Kallenberg and Ledwina, 1995b suggested the linearization as follows
@@
This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\hat{\Psi}^{2}_{\hat{K}}$. $\square$
Implement issues in simulation studies.
In our simulation study, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=0}^{\infty}$. The properties of orthonormal Legendre polynomials lead to the following expressions for constants $c_{1k}$ and $c_{2k}$:
@@
Additional simulation results.
The results of Experiment III are reported in Tables 14, 15, 16 and Figure 5. The main findings are broadly consistent with those in Experiment III, with two additional observations. First, Table 14 indicates that the limiting $\chi^{2}_{1}$ distribution for the data-driven test yields valid control of the Type I error. Second, the empirical distribution of $\hat{K}$ exhibits increased stability under both the null and the alternative hypotheses, with instances where it degenerates to a single value with probability one. Both phenomena are due to the increase in total sample size resulting from the larger number of groups, which is consistent with the theoretical results.
Table 14: Empirical rejection rates under $H_{0}$ in Experiment III, for fixed $K$ ($K=1,\ldots,5$), the data-driven test with limiting $\chi^{2}_{1}$ null distribution ($\hat{K}\&\chi^{2}_{1}$), and the data-driven test with approximated null distribution $H(x)$ ($\hat{K}\&H(x)$). The empirical power under $H_{1}$ equals 1 across all settings.
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\hat{K}\&\chi^{2}_{1}$ $\hat{K}\&H(x)$
10 0.028 0.046 0.062 0.036 0.054 0.060 0.048
20 0.046 0.066 0.078 0.052 0.068 0.056 0.052
30 0.050 0.056 0.044 0.040 0.048 0.044 0.042
40 0.054 0.054 0.042 0.040 0.044 0.058 0.052
50 0.052 0.062 0.052 0.056 0.054 0.048 0.042
60 0.062 0.050 0.060 0.048 0.048 0.052 0.046
70 0.068 0.052 0.052 0.054 0.044 0.058 0.054
80 0.052 0.034 0.044 0.062 0.052 0.048 0.046
90 0.048 0.048 0.034 0.050 0.058 0.042 0.040
100 0.064 0.048 0.052 0.048 0.052 0.058 0.054
110 0.044 0.048 0.058 0.058 0.032 0.060 0.056
120 0.052 0.058 0.066 0.044 0.052 0.080 0.078
130 0.046 0.050 0.038 0.058 0.062 0.058 0.052
140 0.058 0.046 0.032 0.054 0.046 0.056 0.050
150 0.046 0.040 0.060 0.050 0.054 0.038 0.038
Table 15: Empirical frequency of $\hat{K}$ under $H_{0}$ in Experiment III
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0.994 0.006 0 0 0
20 1 0 0 0 0
30 0.992 0.008 0 0 0
40 0.994 0.006 0 0 0
50 0.988 0.012 0 0 0
60 0.994 0.006 0 0 0
70 0.998 0.002 0 0 0
80 0.992 0.008 0 0 0
90 0.996 0.004 0 0 0
100 0.994 0.006 0 0 0
110 0.996 0.004 0 0 0
120 0.998 0.002 0 0 0
130 0.990 0.010 0 0 0
140 0.994 0.006 0 0 0
150 0.998 0.002 0 0 0
Table 16: Empirical frequency of $\hat{K}$ under $H_{1}$ in Experiment III
$m$ $\hat{K}=1$ $\hat{K}=2$ $\hat{K}=3$ $\hat{K}=4$ $\hat{K}=5$
10 0 0 0 0.120 0.880
20 0 0 0 0.024 0.976
30 0 0 0 0 1
40 0 0 0 0 1
50 0 0 0 0 1
60 0 0 0 0 1
70 0 0 0 0 1
80 0 0 0 0 1
90 0 0 0 0 1
100 0 0 0 0 1
110 0 0 0 0 1
120 0 0 0 0 1
130 0 0 0 0 1
140 0 0 0 0 1
150 0 0 0 0 1
Figure 5: The sample means and error bars of $\hat{K}$ in Experiment III

### v4

@@ A.7 Proof of Theorem 10
from the asymptotic $\chi_{1}^{2}$ distribution and asymptotic independence of $N\widehat{\Psi}^{2}_{1}$, $N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})$. For $x\geq 2\log N$, we have
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,\widehat{K}=2\right)$
$\approx$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
$=$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)-\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right).$
As $N(\widehat{\Psi}^{2}_{2}-\widehat{\Psi}^{2}_{1})$ is approximately $\chi_{1}^{2}$ distributed, we obtain
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
@@
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
$\leq$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>x\right)$
$\leq$ $\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}>2\log N\right)$
@@
where we use the distribution function of the limiting $\chi^{2}_{2}$ for $N\widehat{\Psi}^{2}_{2}$. Hence $\mathbb{P}(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N)$ converges to 0 much slower than $\mathbb{P}(N\widehat{\Psi}^{2}_{2}>x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N)$, and thus the latter can be neglected. Therefore, for $x\geq 2\log N$, we have the approximation
$\mathbb{P}\left(N\widehat{\Psi}^{2}_{1}\leq x,N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\leq\log N\right)+\mathbb{P}\left(N\widehat{\Psi}^{2}_{2}-N\widehat{\Psi}^{2}_{1}\geq\log N\right)$
@@
For $\log N<x<2\log N$, kallenberg1995data suggested the linearization as follows
@@
This completes the derivation of the approximation (4.3) for the null distribution of the test statistic $N\widehat{\Psi}^{2}_{\widehat{K}}$. $\square$
A.9 Implement issues in simulation studies
In our simulation study, for the proposed smooth tests, we adopt the orthonormal Legendre polynomials on $[0,1]$ for $\{\pi_{k}\}_{k=1}^{\infty}$. The properties of orthonormal Legendre polynomials lead to the following expressions for constants $c_{1k}$ and $c_{2k}$:
@@ A.9 Implement issues in simulation studies
A.10 Additional simulation results
The results for Experiment III are reported in Tables 13, 14 and Figure 5. They are broadly consistent with those obtained in Experiment III, with several additional findings. First, when $J=10$ and $m$ is relatively large, the total sample size $N$ exceeds the applicability range of the Shapiro–Wilk test (which requires $N\leq 3000$), and hence the corresponding results are not reported in Table 13. In contrast, our proposed test is not subject to this limitation and remains applicable. Second, Table 13 shows that the limiting $\chi^{2}_{1}$ distribution for the data-driven test provides valid control of the Type I error. Third, the empirical distribution of $\widehat{K}$ becomes more stable under both the null and alternative hypotheses, in some cases degenerating to a single value with probability one. These observations can be attributed to the increase in total sample size induced by a larger number of groups, which is in line with the theoretical results.
Table 13: Empirical rejection rates under $H_{0}$ in Experiment III. The empirical power under $H_{1}$ equals 1 for all sample sizes and all considered tests.
Smooth tests JB KS
$m$ $K=1$ $K=2$ $K=3$ $K=4$ $K=5$ $\widehat{K}\&\chi^{2}_{1}$ $\widehat{K}\&H(x)$
10 0.028 0.046 0.062 0.036 0.054 0.060 0.048 0.050 0
20 0.046 0.066 0.078 0.052 0.068 0.056 0.052 0.040 0
30 0.050 0.056 0.044 0.040 0.048 0.044 0.042 0.054 0
40 0.054 0.054 0.042 0.040 0.044 0.058 0.052 0.046 0
50 0.052 0.062 0.052 0.056 0.054 0.048 0.042 0.066 0
60 0.062 0.050 0.060 0.048 0.048 0.052 0.046 0.050 0
70 0.068 0.052 0.052 0.054 0.044 0.058 0.054 0.040 0
80 0.052 0.034 0.044 0.062 0.052 0.048 0.046 0.042 0
90 0.048 0.048 0.034 0.050 0.058 0.042 0.040 0.044 0
100 0.064 0.048 0.052 0.048 0.052 0.058 0.054 0.050 0
110 0.044 0.048 0.058 0.058 0.032 0.060 0.056 0.048 0.002
120 0.052 0.058 0.066 0.044 0.052 0.080 0.078 0.046 0
130 0.046 0.050 0.038 0.058 0.062 0.058 0.052 0.046 0
140 0.058 0.046 0.032 0.054 0.046 0.056 0.050 0.056 0
150 0.046 0.040 0.060 0.050 0.054 0.038 0.038 0.042 0.002
Table 14: Empirical frequency of $\widehat{K}$ in Experiment III
Under $H_{0}$ Under $H_{1}$
$m$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$ $\widehat{K}=1$ $\widehat{K}=2$ $\widehat{K}=3$ $\widehat{K}=4$ $\widehat{K}=5$
10 0.994 0.006 0 0 0 0 0 0 0.120 0.880
20 1 0 0 0 0 0 0 0 0.024 0.976
30 0.992 0.008 0 0 0 0 0 0 0 1
40 0.994 0.006 0 0 0 0 0 0 0 1
50 0.988 0.012 0 0 0 0 0 0 0 1
60 0.994 0.006 0 0 0 0 0 0 0 1
70 0.998 0.002 0 0 0 0 0 0 0 1
80 0.992 0.008 0 0 0 0 0 0 0 1
90 0.996 0.004 0 0 0 0 0 0 0 1
100 0.994 0.006 0 0 0 0 0 0 0 1
110 0.996 0.004 0 0 0 0 0 0 0 1
120 0.998 0.002 0 0 0 0 0 0 0 1
130 0.990 0.010 0 0 0 0 0 0 0 1
140 0.994 0.006 0 0 0 0 0 0 0 1
150 0.998 0.002 0 0 0 0 0 0 0 1
Figure 5: The sample means and error bars of $\widehat{K}$ in Experiment III